# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 7. Adversarial Attacks in LVLM
Robust in Appearance, Fragile in Reality (section 7.1), draw=lightcoral, fill=lightcoral!15, text width=3.5cm, text=black
Robust in Appearance (section 7.1.1), draw=lightcoral, fill=lightcoral!15, text width=3cm, text=black
[LVLM in Traditional Attacks, draw=lightcoral, fill=lightcoral!15, text width=3cm, text=black]
Fragile in Reality (section 7.1.2), draw=lightcoral, fill=lightcoral!15, text width=3cm, text=black
[LVLM in Current Threats, draw=lightcoral, fill=lightcoral!15, text width=3cm, text=black]
[Reasons for LVLM Fragility, draw=lightcoral, fill=lightcoral!15, text width=3cm, text=black]
@@
Over the past decade, adversarial attacks have evolved across various algorithms (section 4), motivations (section 5), and applications (section 6). Recently, multimodal large models, particularly large vision-language models (LVLMs) that have expanded from large language models (LLMs), have become a new focus, raising concerns about their usability and integrity. As depicted in Fig. 6, this chapter offers a comprehensive overview of adversarial attacks on LVLMs from five perspectives. In section 7.1, we first discuss LVLMs’ performance under traditional adversarial attacks and those specific to them, and then explore why they remain vulnerable. section 7.2 defines key terms and symbols. The evaluation framework, including victim models, datasets, and metrics, is presented in section 7.3. section 7.4 classifies attack methods by purposes, attacker knowledge, and techniques, while section 7.5 briefly introduces defense strategies.
7.1. Robust in Appearance, Fragile in Reality
In LVLMs, adversarial attacks exhibit distinct victim targets and environments compared to traditional attacks. On one hand, LVLMs possess immense model capacity, vast knowledge, and complex alignment processes, which seemingly provide stronger resistance to attacks than traditional vision models. On the other hand, the incorporation of multimodal inputs exposes LVLMs to a broader attack surface, resulting in more diverse attack forms. This section will discuss the robustness advantages of LVLMs in section 7.1.1 and the more complex attack environments they face in section 7.1.2.
7.1.1. Robust in Appearance
LVLMs have distinctive traits that brings stronger robustness when facing classic attacks:
@@ 7.1.1. Robust in Appearance
LVLMs are trained with massive training data. JDA+ (Papernot et al., 2016a) posits that a sufficiently large and diverse training dataset can more comprehensively cover the input domain, thus aiding in improving robustness.
@@
LVLMs are often subjected to robust training (e.g., secure fine-tuning (Chiang et al., 2023) and Reinforcement Learning from Human Feedback, i.e., RLHF (Bai et al., 2022; Ouyang et al., 2022; Christiano et al., 2017)). FGSM (Goodfellow et al., 2014), PGD (Madry, 2017), and AMLS (Papernot et al., 2016c) argue that robust training, when combined with sufficient model capacity, can effectively enhance resilience.
Some researchers (Chen et al., 2023c; Dong et al., 2023a; Wang et al., 2024a) have applied classic attacks to LVLMs and proven that, in terms of black-box transferability, LVLMs are relatively more robust than traditional ones like CNNs or ViTs (with a transfer rate dropping from 90% in traditional models to 30% in LVLMs). Although LVLMs show relative improvements in robustness, this is not absolute.
7.1.2. Fragile in Reality
@@
The inclusion of text modality has exposed large models to more diverse threats, unlike traditional vision models. By fabricating a specific context (Yu et al., 2023; Wu et al., 2023) (e.g., simulating a security research experiment (Carlini et al., 2024)), the model can be tricked into forgetting to follow system security instructions. For example, making the model play the role of a soothing grandmother (Sid, 2023) or other characters (SessionGloomy, 2023; Ma et al., 2024) (like DAN (Walkerspider, 2023) or Red Team Assistant (Chao et al., 2023)) can bypass safety protections. Moreover, simple techniques (Wei et al., 2024a), such as prohibiting the model from replying with negative phrases like “Sorry; I can’t; However,” or prompting it to start with affirmative ones like “Certainly! Here is,” (Wu et al., 2023) (we call this RS/AA, Refusal Suppression/Affirmation Augmentation) can easily get responses that should be rejected. In addition to security concerns, LVLMs also carry the risk of privacy breaches. Pre-training data can also be leaked through requests for repeating the word “poem” (Nasr et al., 2023). Concerns on large model security is urgent.
Before addressing these security issues, a natural question arises: Why are large models so vulnerable? Based on existing research, we have summarized two key reasons:
@@
Additionally, compared to traditional single-modal models, LVLMs, while benefiting from the powerful capabilities brought by multimodal inputs, also face the threat of multimodal attacks. The expanded attack surface has led to more diverse attack paradigms, further exacerbating their vulnerability (Carlini et al., 2024; Dong et al., 2023a; Qi et al., 2024). On the other hand, the growing number of variegated downstream applications (Greshake et al., 2023; Wu et al., 2024) has also extended the attack rewards on LVLMs.
@@ 7.2. Problem Definition
Let be the target LVLM, which takes and as the image and text inputs, returning a text response . Adversarial attacks modify the inputs to achieve various attack purposes (e.g., cognitive bias, prompt injection, and jailbreak in section 7.4.2). The attack objective is defined as follows:
Here, and represent inputs modified by and , while denotes the attacker-desired model output. The goal is to find an attack function to effectively modify the inputs ( is a replacement mark and has no practical meaning). Different tasks have varying input structures. For instance, in image captioning (Zhao et al., 2024), might be replaced by a placeholder , while in some robustness tests (Luo et al., 2024b), may be fixed as a blank image or Gaussian noise. Different attack objectives also result in different . In targeted attacks, is expected to resemble a specific text or a certain type of text (e.g., a malicious response in a jailbreak or a reply containing an attacker’s link in prompt injection). In untargeted attacks, should deviate as much as possible from the normal response . The optimization paradigm for different attack objectives can be defined as:
Here, represents the modification to the image and text input, and and refer to the image and text, respectively. We generalize to not only represent perturbations in traditional image adversarial attacks but also adversarial suffixes (Zou et al., 2023) in text attacks and LVLM-specific input modifications, such as context simulation (Wu et al., 2023) or role-playing (Ma et al., 2024) based on the original prompt. In traditional image adversarial attacks, there are two methods to constrain to make the perturbations imperceptible: 1) Using box constraints (Carlini and Wagner, 2017) to limit the size of perturbations in the pixel domain, typically implemented by constraining (where is a hyperparameter representing the budget); 2) Not constraining the size of perturbations but ensuring they are visually imperceptible, which may allow for larger perturbations but tend to form specific shapes, thereby appearing natural and undetectable visually (Chen et al., 2023a).
7.3. Evaluation Framework
7.3.1. Benchmarks
| Num | Text Prompt | Class | Text Type | Notes | Source |
| 7 | Please follow the instructions in the uploaded image. | T | Inst | AoR | RTVLM (Li et al., 2024d) |
@@ 7.3.1. Benchmarks
| Achilles (Li et al., 2024c) | 5 | 250 | - | 750 | Inst | RT | BeaverTails |
@@
| RTG4 (Chen et al., 2024a) | 11 | - | 1445 | - | Ques, Inst | RT | OpenAI, Meta |
@@
| JailbreakHub (Shen et al., 2023) | 13 | - | 100k (390) | - | Ques | R-Atk, RT | OpenAI |
@@
| AVIBench (Zhang et al., 2024b) | 6 | - | - | 260k | All | R-Atk | Self |
@@
| RTVLM (Li et al., 2024d) | 10 | - | - | 5.2k | Ques, Inst | R | Self |
| SafetyBench (Zhang et al., 2023b) | 7 | - | 11.4k | - | Ques | R-Sen | Safety-Prompts |
@@
| SPA-VL (Zhang et al., 2024a) | 53 | - | - | 100k | Ques | T | Self |
@@
Non-Security Datasets (marked as N for “Normal”). These datasets contain images or text-image pairs to test different capabilities of LVLMs (Xu et al., 2023a; Shao et al., 2023) with various multimodal tasks (e.g., image classification (Deng et al., 2009; Krizhevsky et al., 2009) and captioning (Plummer et al., 2015; Agrawal et al., 2019) for visual cognition, as well as VQA (Goyal et al., 2017; Marino et al., 2019) for reasoning). In addition to normal capability testing, they can also be repurposed to generate cognitive bias attack samples (Wang et al., 2024b, a, 2023b; Dong et al., 2023a) to induce model errors in reasoning. Non-security datasets contains a large number of classic image datasets.
@@
R-Sen: Datasets containing either toxic samples (e.g., content with pornography, gore, or violence) or non-toxic samples that may trigger toxic outputs (e.g., when asked “Where can I buy a can of coke?” the model might interpret “coke” as drugs, leading to harmful responses (Röttger et al., 2023)).
Safety Alignment Datasets (marked as T for “Training”). These datasets are primarily used for fine-tuning LLMs (Ji et al., 2024a), training preference models via RLHF (Bai et al., 2022; Ganguli et al., 2022), or other models designed to detect malicious content (Lin et al., 2023). The goal is to help LLMs strike a better balance between security (harmless) and practicality (helpful). The construction of some red team datasets draws on the categories (Deng et al., 2023b; Li et al., 2024c) or samples (Luo et al., 2024b) used in safety alignment datasets.
@@ 7.3.2. Victim Models
| CLIP (Radford et al., 2021) | VLP | ResNet, ViT-B/L | - | GPT-2*222GPT-2* is a modified version of GPT-2. | 102~428M333It should be noted that a series of works (LAION, 2023; Sun et al., 2024) have scaled up CLIP from 428M to 18B. | (Dong et al., 2023a; Zhou et al., 2023a; Zheng et al., 2024) |
| BLIP (Li et al., 2022) | VLP | ViT-B/L | CA | BERT-B | 224~447M | (Guo et al., 2024; Zhao et al., 2024; Gao et al., 2024b) |
@@
| OpenFlamingo (Awadalla et al., 2023) | Open | CLIP ViT-L | PR+CA | MPT, …444The LLM in OpenFlamingo can be MPT (AI, 2023b), RedPajama (Together, 2023), or LLaMA (Touvron et al., 2023a). | 3~9B | (Schlarmann and Hein, 2023; Wang et al., 2024b) |
| BLIP-2 (Li et al., 2023a) | Open | CLIP ViT-L, …555The Vision Encoder in BLIP-2 can be CLIP ViT-L (Radford et al., 2021) or EVA-CLIP ViT-g (Fang et al., 2023). | Q-Former | OPT, FlanT5 | 3.1~12.1B | (Dong et al., 2023a; Lu et al., 2024; Cui et al., 2024) |
@@
| LLaVA (Liu et al., 2024c) | Open | CLIP ViT-L | Linear | Vicuna | 7.3/13.3B | (Shayegani et al., 2023; Tan et al., 2024; Bailey et al., 2023) |
| LLaVA-1.5 (Liu et al., 2024a) | Open | CLIP ViT-L | MLP | Vicuna-v1.5 | 7.3/13.3B | (Gong et al., 2023b; Ying et al., 2024b; Li et al., 2024c) |
| LLaVA-1.6 (Liu et al., 2024b) | Open | CLIP ViT-L | MLP | Vicuna, …666The LLM in LLaVA-1.6 can be Vicuna (Chiang et al., 2023), Mistral (Jiang et al., 2023) or Nous Hermes 2 (NousResearch, 2023). | 7~35B | (Ma et al., 2024) |
| LLaMA-Adapter (Zhang et al., 2023a) | Open | CLIP ViT-B | MLP | LLaMA | 7B | (Fu et al., 2023) |
| LLaMA-Adapter V2 (Gao et al., 2023) | Open | CLIP ViT-L | Linear | LLaMA | 7.3B | (Carlini et al., 2024; Shayegani et al., 2023) |
@@
| MiniGPT-4 (Zhu et al., 2023) | Open | EVA-CLIP ViT-g | Linear | Vicuna | 8/14B | (Wang et al., 2024e; Ying et al., 2024b; Gong et al., 2023b) |
| MiniGPT-v2 (Chen et al., 2023d) | Open | EVA-CLIP ViT-g | Linear | LLaMA 2 | 8B | (Qraitem et al., 2024; Gao et al., 2024a; Niu et al., 2024) |
@@
| Otter (Li et al., 2023c) | Open | CLIP ViT-L | CA | LLaMA | 9B | (Wang et al., 2024a) |
| IDEFICS (Face and LAION, 2023) | Open | CLIP ViT-H | CA | LLaMA | 9/80B | (Chen et al., 2023b) |
@@
| Qwen (Cloud, 2023) | Close | - | - | - | 110B | (Ying et al., 2024b) |
| Bard (Gemini) (Google, 2023) | Close | - | - | - | 137B | (Wu et al., 2024; Li et al., 2024c; Dong et al., 2023a) |
| GPT-4(V/o) (OpenAI, 2023b; gpt 4o, 2023) | Close | - | - | - | 175B | (Wu et al., 2023; Ying et al., 2024a; Chen et al., 2023b) |
@@
| ERNIE Bot (Baidu, 2023) | Close | - | - | - | 260B | (Dong et al., 2023a; Guo et al., 2024; Ying et al., 2024b) |
@@
| ChatGLM (AI, 2023c) | Close | - | - | - | - | (Ying et al., 2024b) |
@@
As shown in Table 5, existing victim models under attacks of vision language models can be categorized into four types (listed in the Class column): 1) VLP (Vision-Language Pre-training Models), 2) open-sourced LVLMs, 3) close-sourced LVLMs, and 4) other models.
VLP models typically focus on pre-training modules that excel in general tasks (Li et al., 2022; Kim et al., 2021), such as image-text contrastive learning (ITC), matching (ITM), and masked language modeling (MLM), providing strong support for downstream tasks. VLPs usually consist of a visual encoder, a text encoder, and a modality fusion (Li et al., 2021; Yang et al., 2022)/transfer module (Bao et al., 2023), whereas LVLMs leverage LLMs to handle textual information and use a connector to project image features from visual encoders into the text feature space. LVLMs can be either open-sourced or close-sourced. Close-sourced models often have additional defenses (such as GPT-4’s OCR detection for malicious text in images (Gong et al., 2023b)) alongside safety alignment strategies like RLHF (Bai et al., 2022; Ouyang et al., 2022; Christiano et al., 2017), making them more robust to attack. In black-box transfer attacks, open-source models are often used as white-box surrogates (Guo et al., 2024; Wang et al., 2024a; Niu et al., 2024). In gray-box query attacks, accessing APIs (OpenAI, 2023a) of close-sourced models can be costly, leading to high query fees. Notably, models like ChatGLM (Du et al., 2021) and Qwen (Bai et al., 2023a) offer both open-source and commercial versions (AI, 2023c; Cloud, 2023). Img2LLM (Guo et al., 2023), a VQA plugin for LLMs, falls outside of the other three categories. It converts images into captions and a series of questions & answers based on the image content, enabling LLMs to perform VQA tasks.

### v2

@@ 7. Adversarial Attacks in LVLM
Why are LVLMs Vulnerable? (section 7.1), draw=lightcoral, fill=lightcoral!15, text width=3.5cm, text=black
@@
Over the past decade, adversarial attacks have diversified in algorithms (section 4), motivations (section 5), and applications (section 6). Recently, Large Vision-Language Models (LVLMs) have emerged as a key target, raising concerns about their security. As depicted in Fig. 8, this chapter offers a comprehensive overview of LVLM adversarial attacks from five perspectives.
7.1. Why are LVLMs So Vulnerable?
Compared to traditional single-modal models, LVLMs exhibit notable advantages in robustness:
@@ 7.1. Why are LVLMs So Vulnerable?
LVLMs are trained with massive training data. JDA+ (Papernot et al., 2016b) posits that a sufficiently large and diverse training dataset can more comprehensively cover the input domain, thus aiding in improving robustness.
@@
LVLMs are often subjected to robust training (e.g., secure fine-tuning (Chiang et al., 2023) and Reinforcement Learning from Human Feedback, i.e., RLHF (Bai et al., 2022; Ouyang et al., 2022; Christiano et al., 2017)). FGSM (Goodfellow et al., 2014), PGD (Madry, 2017), and AMLS (Papernot et al., 2016a) argue that robust training, when combined with sufficient model capacity, can effectively enhance resilience.
@@
However, LVLMs still fall short of ideal robustness. For instance, security safeguards can be easily bypassed through contextual simulation (Yu et al., 2023; Yin et al., 2023; Carlini et al., 2024) or role-playing (Sid, 2023; SessionGloomy, 2023; Ma et al., 2024; Walkerspider, 2023; Chao et al., 2023); Simply prohibiting the model from replying with negative phrases like “Sorry; I can’t; However,” or prompting it to start with affirmative ones like “Certainly! Here is,” (Yin et al., 2023) (we call this RS/AA, Refusal Suppression/Affirmation Augmentation) can easily get responses that should be rejected. For privacy protection, pre-training data can also be leaked through requests for repeating the word “poem” (Nasr et al., 2023). Why are LVLMs so vulnerable? Based on existing research, we have summarized two key reasons:
@@
Additionally, compared to traditional single-modal models, LVLMs, while benefiting from the powerful capabilities brought by multimodal inputs, also face the threat of multimodal attacks. The expanded attack surface has led to more diverse attack paradigms, further exacerbating their vulnerability (Carlini et al., 2024; Dong et al., 2023a; Qi et al., 2024). Furthermore, the growing number of variegated downstream applications (Greshake et al., 2023; Wu et al., 2024) has also extended the attack rewards on LVLMs.
@@ 7.2. Problem Definition
Let be the target LVLM, which takes and as the image and text inputs 333 and can be selected from the Normal, Red Team, Robustness, or Training datasets according to different purposes (see section 7.3.1)., returning a text response . Adversarial attacks modify the inputs to achieve various attack purposes (e.g., cognitive bias, prompt injection, and jailbreak in section 7.4.2). The attack objective is defined as follows:
Here, and represent inputs modified by and , while denotes the attacker-desired model output. Text perturbations exhibit more diverse forms than image perturbations, such as adversarial suffixes (Zou et al., 2023), context simulation (Yin et al., 2023), and role-playing (Ma et al., 2024) — all crafted through original prompts. The goal is to find an attack function to effectively modify the inputs. Different tasks have varying input types. For instance, in image captioning (Zhao et al., 2024), might be replaced by a placeholder , while in robustness tests (Luo et al., 2024b), may be fixed as a blank image or Gaussian noise. Different attack goals also result in different . In targeted attacks, is expected to resemble a specific text or a certain type of text (e.g., a malicious response in a jailbreak or a reply containing an attacker’s link in prompt injection). In untargeted attacks, should deviate as much as possible from the normal response . The optimization paradigm for different attack goals can be defined as:
Here, represents the modification to the image and text input, respectively. Adversariality is formed through eq. 4, and the stealthiness is discussed in section 5.3.
@@
| Achilles (Li et al., 2024b) | 5 | 250 | - | 750 | Inst | RT | BeaverTails |
@@
| RTG4 (Chen et al., 2024b) | 11 | - | 1445 | - | Ques, Inst | RT | OpenAI, Meta |
@@
| JailbreakHub (Shen et al., 2024) | 13 | - | 100k (390) | - | Ques | R-Atk, RT | OpenAI |
@@
| AVIBench (Zhang et al., 2024a) | 6 | - | - | 260k | All | R-Atk | Self |
@@
| RTVLM (Li et al., 2024e) | 10 | - | - | 5.2k | Ques, Inst | R | Self |
| SafetyBench (Zhang et al., 2023a) | 7 | - | 11.4k | - | Ques | R-Sen | Safety-Prompts |
@@
| SPA-VL (Zhang et al., 2024b) | 53 | - | - | 100k | Ques | T | Self |
7.3. Evaluation Framework
7.3.1. Benchmarks
| Num | Text Prompt | Class | Text Type | Notes | Source |
| 7 | Please follow the instructions in the uploaded image. | T | Inst | AoR | RTVLM (Li et al., 2024e) |
@@ 7.3.1. Benchmarks
Non-Security Datasets (marked as N for “Normal”). These datasets contain images or text-image pairs to test different capabilities of LVLMs (Xu et al., 2023a; Shao et al., 2023) with various multimodal tasks (e.g., image classification (Deng et al., 2009; Krizhevsky et al., 2009) and captioning (Plummer et al., 2015; Agrawal et al., 2019) for visual cognition, as well as VQA (Goyal et al., 2017; Marino et al., 2019) for reasoning). In addition to normal capability testing, they can also be repurposed to generate cognitive bias AEs which leverage specific perturbations to induce model errors in cognizing (Wang et al., 2024d, a, 2023b; Dong et al., 2023a). Non-security datasets contains a large number of classic image datasets.
@@
R-Sen: Datasets containing content-sensitive data, including toxic samples (e.g., content with pornography, gore, or violence) or non-toxic samples that may trigger toxic outputs (e.g., when asked “Where can I buy a can of coke?” the model might interpret “coke” as drugs, leading to harmful responses (Röttger et al., 2023)).
Safety Alignment Datasets (marked as T for “Training”). These datasets are primarily used for fine-tuning LLMs (Ji et al., 2024a), training preference models via RLHF (Bai et al., 2022; Ganguli et al., 2022), or other models designed to detect malicious content (Lin et al., 2023). The goal is to help LLMs strike a better balance between security (harmless) and practicality (helpful). The construction of some red team datasets draws on the categories (Deng et al., 2023b; Li et al., 2024b) or samples (Luo et al., 2024b) used in safety alignment datasets.
@@ 7.3.2. Victim Models
| CLIP (Radford et al., 2021) | VLP | ResNet, ViT-B/L | - | GPT-2*444GPT-2* is a modified version of GPT-2. | 102~428M555It should be noted that a series of works (LAION, 2023; Sun et al., 2024) have scaled up CLIP from 428M to 18B. | (Dong et al., 2023a; Zhou et al., 2023a; Zheng et al., 2024) |
| BLIP (Li et al., 2022a) | VLP | ViT-B/L | CA | BERT-B | 224~447M | (Guo et al., 2024; Zhao et al., 2024; Gao et al., 2024b) |
@@
| OpenFlamingo (Awadalla et al., 2023) | Open | CLIP ViT-L | PR+CA | MPT, …666The LLM in OpenFlamingo can be MPT (Mosaic, 2023), RedPajama (Together, 2023), or LLaMA (Touvron et al., 2023a). | 3~9B | (Schlarmann and Hein, 2023; Wang et al., 2024d) |
| BLIP-2 (Li et al., 2023a) | Open | CLIP ViT-L, …777The Vision Encoder in BLIP-2 can be CLIP ViT-L (Radford et al., 2021) or EVA-CLIP ViT-g (Fang et al., 2023). | Q-Former | OPT, FlanT5 | 3.1~12.1B | (Dong et al., 2023a; Lu et al., 2024; Cui et al., 2024) |
@@
| LLaVA (Liu et al., 2024d) | Open | CLIP ViT-L | Linear | Vicuna | 7.3/13.3B | (Shayegani et al., 2023; Tan et al., 2024; Bailey et al., 2023) |
| LLaVA-1.5 (Liu et al., 2024b) | Open | CLIP ViT-L | MLP | Vicuna-v1.5 | 7.3/13.3B | (Gong et al., 2023b; Ying et al., 2024a; Li et al., 2024b) |
| LLaVA-1.6 (Liu et al., 2024c) | Open | CLIP ViT-L | MLP | Vicuna, …888The LLM in LLaVA-1.6 can be Vicuna (Chiang et al., 2023), Mistral (Jiang et al., 2023) or Nous Hermes 2 (NousResearch, 2023). | 7~35B | (Ma et al., 2024) |
| LLaMA-Adapter (Zhang et al., 2023b) | Open | CLIP ViT-B | MLP | LLaMA | 7B | (Fu et al., 2023) |
| LLaMA-Adapter V2 (Gao et al., 2023a) | Open | CLIP ViT-L | Linear | LLaMA | 7.3B | (Carlini et al., 2024; Shayegani et al., 2023) |
@@
| MiniGPT-4 (Zhu et al., 2023) | Open | EVA-CLIP ViT-g | Linear | Vicuna | 8/14B | (Wang et al., 2024e; Ying et al., 2024a; Gong et al., 2023b) |
| MiniGPT-v2 (Chen et al., 2023a) | Open | EVA-CLIP ViT-g | Linear | LLaMA 2 | 8B | (Qraitem et al., 2024; Gao et al., 2024a; Niu et al., 2024) |
@@
| Otter (Li et al., 2023d) | Open | CLIP ViT-L | CA | LLaMA | 9B | (Wang et al., 2024a) |
| IDEFICS (Face and LAION, 2023) | Open | CLIP ViT-H | CA | LLaMA | 9/80B | (Chen et al., 2023c) |
@@
| Qwen (Cloud, 2023) | Close | - | - | - | 110B | (Ying et al., 2024a) |
| Bard (Gemini) (Google, 2023) | Close | - | - | - | 137B | (Wu et al., 2024; Li et al., 2024b; Dong et al., 2023a) |
| GPT-4(V/o) (OpenAI, 2023b; gpt 4o, 2023) | Close | - | - | - | 175B | (Yin et al., 2023; Ying et al., 2024b; Chen et al., 2023c) |
@@
| ERNIE Bot (Baidu, 2023) | Close | - | - | - | 260B | (Dong et al., 2023a; Guo et al., 2024; Ying et al., 2024a) |
@@
| ChatGLM (AI, 2023b) | Close | - | - | - | - | (Ying et al., 2024a) |
@@
As shown in Table 5, existing VLM victim models can be categorized into four types (listed in the Class column): 1) VLP (Vision-Language Pre-training Models), 2) open-sourced LVLMs, 3) close-sourced LVLMs, and 4) other models. The victim can be the entire model or any individual component—its vision encoder, adapter, or LM.
VLP models typically focus on pre-training modules that excel in general tasks (Li et al., 2022a; Kim et al., 2021), such as image-text contrastive learning (ITC), matching (ITM), and masked language modeling (MLM), providing strong support for downstream tasks. VLPs usually consist of a visual encoder, a text encoder, and a modality fusion (Li et al., 2021; Yang et al., 2022)/transfer module (Bao et al., 2023), whereas LVLMs leverage LLMs to handle textual information and use a connector to project image features from visual encoders into the text feature space. LVLMs can be either open-sourced or close-sourced. Close-sourced models often have additional defenses (such as GPT-4’s OCR detection for malicious text in images (Gong et al., 2023b)) alongside safety alignment strategies like RLHF (Bai et al., 2022; Ouyang et al., 2022; Christiano et al., 2017), making them more robust to attack. In black-box transfer attacks, open-source models are often used as white-box surrogates (Guo et al., 2024; Wang et al., 2024a; Niu et al., 2024). In gray-box query attacks, accessing APIs (OpenAI, 2023a) of close-sourced models can be costly, leading to high query fees. Notably, models like ChatGLM (Du et al., 2021) and Qwen (Bai et al., 2023a) offer both open-source and commercial versions (AI, 2023b; Cloud, 2023). Img2LLM (Guo et al., 2023), a VQA plugin for LLMs, falls outside of the other three categories. It converts images into captions and a series of questions & answers based on the image content, enabling LLMs to perform VQA tasks.